# Algebraic Robustness Verification of Neural Networks

### v1

@@ Abstract
We formulate formal robustness verification of neural networks as an algebraic optimization problem. We leverage the Euclidean Distance (ED) degree, which is the generic number of complex critical points of the distance minimization problem to a classifier’s decision boundary, as an architecture-dependent measure of the intrinsic complexity of robustness verification. To make this notion operational, we define the associated ED discriminant, which characterizes input points at which the number of real critical points changes, distinguishing test instances that are easier or harder to verify. We provide an explicit algorithm for computing this discriminant. We further introduce the parameter discriminant of a neural network, identifying parameters where the ED degree drops and the decision boundary exhibits reduced algebraic complexity. We derive closed-form expressions for the ED degree for several classes of neural architectures, as well as formulas for the expected number of real critical points in the infinite-width limit. Finally, we present an exact robustness certification algorithm based on numerical homotopy continuation, establishing a concrete link between metric algebraic geometry and neural network verification.
Keywords: ED degree, ED discriminant, Kac-Rice formula, parameter discriminant, robustness verification, decision boundary
@@ 1 Introduction
The deployment of neural networks in safety-critical settings has led to a growing demand for formal robustness guarantees, ensuring that small perturbations of the input do not lead to incorrect or unsafe predictions. Robustness verification, certifying that a classifier’s output remains invariant over a specified neighborhood of an input, is therefore a central problem in modern machine learning.
A wide range of verification methods have been proposed, including convex relaxations (Salman et al., 2019), mixed-integer programming (Tjeng et al., 2019), satisfiability modulo theories (Katz et al., 2017), and abstract interpretation (Gehr et al., 2018). While these approaches have achieved success in various settings (Weng et al., 2018; Zhang et al., 2018; Botoeva et al., 2020; Brix et al., 2023), they all face a tradeoff between exactness and scalability: exact methods often suffer from exponential complexity or strong dependence on specific architectures (Meng et al., 2022; Liu et al., 2024; Froese et al., 2025), whereas relaxed methods can achieve fast computation, but they may be loose (Singh et al., 2018; Wong and Kolter, 2018; Salman et al., 2019) and incapable to verify difficult instances. There is limited theoretical understanding of how the intrinsic structure of a neural network governs the complexity of robustness verification (Ferlez and Shoukry, 2021; König et al., 2024).
We propose an algebraic framework for robustness verification, treating neural networks and their decision boundaries as algebraic objects. For polynomial classifiers and, more generally, decision boundaries that are algebraic or semi-algebraic, robustness verification can be formulated as an algebraic distance minimization problem (see Fig.˜1). In this case, the decision boundary, which is a set of input points satisfying an equation of the form $f_{i}(\bm{x})-f_{j}(\bm{x})=0$, forms an algebraic variety. Robustness verification amounts to certifying that the distance from a given test point $\bm{\xi}$ to the decision boundary is bounded below by some $\epsilon$. This perspective naturally connects robustness verification to tools from metric algebraic geometry, where the complexity of distance optimization is captured by algebraic invariants of the underlying variety.
@@
Our central complexity measure is the Euclidean Distance (ED) degree, a fundamental invariant from algebraic geometry, defined as the generic number of complex critical points of the distance function to the decision boundary (Draisma et al., 2016; Maxim et al., 2020; Breiding et al., 2024; Lai et al., 2025). In the context of robustness verification, the ED degree provides an architecture-dependent notion of verification complexity, independent of optimization heuristics. We leverage this notion by introducing two discriminants. First, the ED discriminant characterizes input points at which the number of real critical points changes, thereby identifying test instances that are easier or harder to certify. We provide explicit algorithms for its computation. Second, the parameter discriminant identifies network parameters for which the ED degree drops, indicating reduced algebraic complexity of the decision boundary. We derive closed formulas for the ED degree for several classes of network architectures, as well as for the expected number of real critical points in the infinite-width limit. Finally, we present an exact robustness certification algorithm based on numerical homotopy continuation, yielding provably correct certificates whenever the algebraic problem can be solved.
@@ 1.1 Contributions
  • In Section 2, we formulate robustness verification as a metric projection problem onto algebraic varieties, providing an exact KKT-characterization and a relaxation.
  • In Section 3, we propose the ED degree as a notion of verification complexity and outline its computation (Alg. 1).
  • In Section 3.1, we derive closed-form expressions for the ED degree of decision boundaries under generic parameters for wide and bottleneck architectures (Prop. 3.8 and Prop. 3.6), and give numerical results for intermediate regimes.
  • In Section 3.2, we derive a general formula for the expected number of real critical points (Prop. 3.10), and a result for Neural Network Gaussian Processes (NNGP) (Cor. 3.11).
  • In Section 4.1, we define the ED discriminant, which classifies test instances by the complexity of the verification task. We provide a symbolic algorithm for its computation (Alg. 2), along with examples.
  • In Section 4.2, we introduce the parameter discriminant, which classifies the parameters of a neural network by the verification complexity. We establish a complete characterization for quadric boundaries (Thm. 4.6) and derive the parameter discriminant as a stratified hypersurface.
An implementation of our algorithms is provided at: https://github.com/edwardduanhao/AlgebraicVerification.
@@ 1.2 Related Works
We connect these lines of research by introducing algebraic complexity measures based on the Euclidean Distance degree and associated discriminants for robustness verification. Our framework treats verification difficulty as an intrinsic function of architecture and parameters, and yields exact algorithms grounded in metric algebraic geometry.
@@ 2.1 Robustness Verification via Metric Projection
If $\delta(\epsilon)\geq 0$, then the score of class $c$ exceeds all others in $\mathbb{B}(\bm{\xi},\epsilon)$, thereby certifying robustness at $\bm{\xi}$.
@@
Note that $\mathcal{B}_{c,c^{\prime}}^{\bm{\theta}}\subseteq\mathcal{V}_{c,c^{\prime}}^{\bm{\theta}}$, since the decision boundary requires extra constraints $f_{\bm{\theta},i}(\bm{x})-f_{\bm{\theta},c}(\bm{x})\leq 0$ for all $i\neq c,c^{\prime}$.
We quantify the robustness margin as the minimum distance from the test point $\bm{\xi}$ to the active decision boundary,
@@
The condition $\gamma\geq\epsilon$ implies that the ball $\mathbb{B}(\bm{\xi},\epsilon)$ does not intersect any active decision boundary, i.e.,
@@
Equivalently, for every $\bm{x}\in\mathbb{B}(\bm{\xi},\epsilon)$ and every $c^{\prime}\neq c$, we have $f_{\bm{\theta},c}(\bm{x})\geq f_{\bm{\theta},c^{\prime}}(\bm{x})$, so the pairwise margins remain nonnegative everywhere in the ball, and $\delta(\epsilon)\geq 0$. By continuity of the logits, the predicted class then remains invariant within $\mathbb{B}(\bm{\xi},\epsilon)$, thereby certifying robustness at $\bm{\xi}$.
@@ 2.2 Algebraic Optimization Formulation
We focus on the case where the classifier $f_{\bm{\theta}}$ is algebraic, meaning each component $f_{\bm{\theta},i}$ is a polynomial of the input $\bm{x}$. In this setting, $\mathcal{V}_{c,c^{\prime}}^{\bm{\theta}}$ is a real algebraic variety, as it is the set of zeros of the polynomial $f_{\bm{\theta},c}(\bm{x})-f_{\bm{\theta},c^{\prime}}(\bm{x})$. Imposing the additional inequality constraints $f_{\bm{\theta},i}(\bm{x})-f_{\bm{\theta},c}(\bm{x})\leq 0$ for all $i\neq c,c^{\prime}$ restricts this to a semialgebraic set $\mathcal{B}_{c,c^{\prime}}^{\bm{\theta}}$.
@@
$\mathcal{L}(\bm{x},\lambda,\bm{\mu})$={} $\frac{1}{2}\,d_{\bm{\xi}}(\bm{x})+\lambda\bigl(f_{\bm{\theta},c^{\prime}}(\bm{x})-f_{\bm{\theta},c}(\bm{x})\bigr)$
$+\sum_{i\neq c,c^{\prime}}\mu_{i}\bigl(f_{\bm{\theta},i}(\bm{x})-f_{\bm{\theta},c}(\bm{x})\bigr),$
@@
Since all equations are polynomial, the KKT system can be solved using tools from numerical algebraic geometry. We consider specifically homotopy continuation because it is a complete numerical method guaranteed to find all isolated complex solutions (Li, 1997; Sommese et al., 2001; Sommese and Wampler, 2005; Sommese et al., 2005; Duff et al., 2018). This computes all complex solutions to the stationarity, complementary slackness, and equality constraints. Subsequently, one retains the minimal real solution satisfying the remaining feasibility conditions. Fig.˜2 illustrates the concept, with details in Section˜A.
@@
The proof is provided in Section˜B.1. As a consequence, we may consider the relaxed optimization problem
@@
These yield a square system of $n+1$ equations in $n+1$ unknowns, instead of n+k1n+k-1, resulting in substantial computational savings when the number of classes $k$ is large. We provide an implementation in Section˜C. [^1] Our implementation is available at: https://github.com/edwardduanhao/AlgebraicVerification
@@ 3 ED Degree of a Decision Boundary
The ED degree of $\mathcal{V}_{c,c^{\prime}}^{\bm{\theta}}$ is as an upper bound on the optimization complexity: for our data point $\bm{\xi}$, the number of complex critical points is bounded above by $\mathrm{EDdegree}(\mathcal{V}_{c,c^{\prime}}^{\bm{\theta}})$, with equality holding when $\bm{\xi}$ is generic.

### v2

# Robustness Verification of Polynomial Neural Networks

@@ Abstract
We study robustness verification of neural networks via metric algebraic geometry. For polynomial neural networks, certifying a robustness radius amounts to computing the distance to the algebraic decision boundary. We use the Euclidean distance (ED) degree as an intrinsic measure of the complexity of this problem, analyze the associated ED discriminant, and introduce a parameter discriminant that detects parameter values at which the ED degree drops. We derive formulas for the ED degree for several network architectures and characterize the expected number of real critical points in the infinite-width limit. We develop symbolic elimination methods to compute these quantities and homotopy-continuation methods for exact robustness certification. Finally, experiments on lightning self-attention modules reveal decision boundaries with strictly smaller ED degree than generic cubic hypersurfaces of the same ambient dimension.
Keywords: ED degree, ED discriminant, Kac-Rice formula, parameter discriminant, robustness verification, decision boundary, polynomial networks
MSC classes: 14Q65, 65H14, 14Q30, 90C23, 60G15, 68T07
@@ 1 Introduction
The deployment of artificial neural networks in safety-critical settings has created a growing demand for formal robustness guarantees, ensuring that small perturbations of the input do not lead to incorrect or unsafe predictions. Robustness verification, i.e., certifying that the output prediction remains invariant over a prescribed neighborhood of a given input, has therefore emerged as a central problem in modern applications. In this paper, we study this problem for classifiers implemented as polynomial neural networks and, more generally, for classifiers whose decision boundaries are algebraic or semi-algebraic. We aim to advance theoretical frameworks to better understand how robustness verification is influenced by (i) the architecture of the neural network; (ii) the particular test point; and (iii) the particular parameter of the network.
A wide range of verification methods have been proposed, including convex relaxations (Salman et al., 2019), mixed-integer programming (Tjeng et al., 2019), satisfiability modulo theories (Katz et al., 2017), and abstract interpretation (Gehr et al., 2018). While these approaches have achieved notable success in various settings (Weng et al., 2018; Zhang et al., 2018; Botoeva et al., 2020; Brix et al., 2023), they exhibit a fundamental trade-off between computational tractability and exactness: methods that provide exact guarantees may incur exponential complexity or depend strongly on architectural features (Meng et al., 2022; Liu et al., 2024; Froese et al., 2025), whereas scalable relaxations enable efficient computation but may produce loose certificates and fail on challenging instances (Singh et al., 2018; Wong and Kolter, 2018; Salman et al., 2019). There remains limited theoretical understanding of how the intrinsic structure of a neural network governs the complexity of robustness verification (Ferlez and Shoukry, 2021; König et al., 2024).
We propose an algebraic framework for quantifying the complexity of robustness verification focusing on polynomial models and the intrinsic properties of the decision boundary. The decision boundary of a classifier consists of input points at which the predicted class changes, which are points satisfying equations of the form $f_{i}(\bm{x})-f_{j}(\bm{x})=0$. Robustness verification at a test point $\bm{\xi}$ then amounts to certifying that its distance to the decision boundary is bounded below by some $\epsilon$ (see Fig.˜1). For polynomial networks, this decision boundary is an algebraic variety. This allows us to connect robustness verification to metric algebraic geometry, where the complexity of the associated distance minimization problem is captured by algebraic invariants of the underlying variety.
@@
Our central complexity measure is the Euclidean Distance (ED) degree, a fundamental invariant from algebraic geometry, defined as the generic number of complex critical points of the distance function to an algebraic variety (Draisma et al., 2016; Maxim et al., 2020; Breiding et al., 2024a; Lai et al., 2025). In the context of robustness verification, the ED degree provides an architecture-dependent measure of verification complexity that is independent of optimization heuristics. We further study the associated ED discriminant and introduce a parameter discriminant. The ED discriminant characterizes input points at which the number of real critical points changes, while the parameter discriminant identifies network parameters for which the ED degree drops, indicating a reduction in the algebraic complexity of the decision boundary.
We derive closed-form expressions for the ED degree and characterize the parameter discriminant for several classes of polynomial neural network architectures. We further analyze the expected number of real critical points using the Kac-Rice framework. For shallow networks in the infinite-width limit, we obtain explicit formulas and dimension-dependent bounds in terms of the degree of the activation function. On the computational side, we develop symbolic and numerical algorithms for computing the ED degree and discriminants, and introduce an exact robustness certification algorithm based on numerical homotopy continuation. This yields provably correct certificates whenever the underlying algebraic problem can be solved. These results show that, by analyzing the algebraic invariants of the decision boundary, one can systematically quantify how verification complexity depends on the architecture, the network parameters, and the test point.
@@ 1.1 Contributions
  • In Section 2, we formulate robustness verification as a metric projection problem onto algebraic varieties, providing an exact KKT-characterization and a relaxation. We also give a numerical homotopy-continuation-based algorithm for exact robustness certification (Algorithm 1).
  • In Section 3, we propose the ED degree as a measure of robustness verification complexity and describe its computation (Algorithm 2).
  • In Section 3.1, we derive closed-form expressions for the ED degree of decision boundaries of polynomial neural networks under generic parameters, for both wide and bottleneck architectures (Proposition 3.8 and Proposition 3.6), and complement these with numerical results for intermediate regimes.
  • In Section 3.2, we study the expected real ED degree of the decision boundary of polynomial networks via Kac-Rice methods. We prove a general theorem for the expected number of real critical points (Theorem 3.10). For shallow networks in the infinite-width limit, we obtain an exact formula for one-dimensional input data (Proposition 3.12) and an upper bound for arbitrary input dimensions (Theorem 3.11).
  • In Section 4.1, we study the ED discriminant, which classifies test instances over the input space by the corresponding complexity of the verification task. We provide a symbolic algorithm for its computation (Algorithm 3), together with examples.
  • In Section 4.2, we introduce the parameter discriminant, which classifies the parameters of a neural network by the corresponding complexity of the verification task on generic test data. We establish a complete characterization for quadric boundaries (Theorem 4.6) and derive the parameter discriminant as a stratified hypersurface.
  • In Section 5, we report numerical results on the ED degree and ED discriminant for lightning self-attention modules.
An implementation of algorithms and experiments is provided at: https://github.com/edwardduanhao/AlgebraicVerification.
@@ 1.2 Related Works
Metric algebraic geometry
The Euclidean distance degree and the ED discriminant arise naturally in nearest-point problems on algebraic varieties and are central objects in metric algebraic geometry (Draisma et al., 2016; Maxim et al., 2020; Breiding et al., 2024a; Lai et al., 2025). This literature develops both the geometric theory and the computational methods for Euclidean distance optimization, including the study of critical points, discriminants, and related complexity measures (Breiding et al., 2024b). Our work applies this perspective to robustness verification for neural networks, where the relevant nearest-point problem is projection to the decision boundary.
@@ 2.1 Robustness Verification via Metric Projection
If $\delta(\epsilon)>0$, then the score of class $c$ exceeds all others in $\mathbb{B}(\bm{\xi},\epsilon)$, thereby certifying robustness at $\bm{\xi}$.
@@
Note that $\mathcal{B}_{c,c^{\prime}}^{\bm{\theta}}\subseteq\mathcal{V}_{c,c^{\prime}}^{\bm{\theta}}$, since the decision boundary requires extra constraints $f_{\bm{\theta},i}(\bm{x})-f_{\bm{\theta},c}(\bm{x})\leq 0$ for all $i\neq c,c^{\prime}$. We quantify the robustness margin as the minimum distance from the test point $\bm{\xi}$ to the active decision boundary by defining
@@
The condition $\gamma>\epsilon$ implies that the ball $\mathbb{B}(\bm{\xi},\epsilon)$ does not intersect any active decision boundary, i.e.,
@@
Equivalently, for every $\bm{x}\in\mathbb{B}(\bm{\xi},\epsilon)$ and every $c^{\prime}\neq c$, we have $f_{\bm{\theta},c}(\bm{x})>f_{\bm{\theta},c^{\prime}}(\bm{x})$, so the pairwise margins remain nonnegative everywhere in the ball, and $\delta(\epsilon)>0$. By continuity of the logits, the predicted class then remains invariant within $\mathbb{B}(\bm{\xi},\epsilon)$, thereby certifying robustness at $\bm{\xi}$.
@@ 2.2 Algebraic Optimization Formulation
Suppose each logit $f_{\bm{\theta},i}$ is a polynomial in $\bm{x}$, so $\mathcal{V}_{c,c^{\prime}}^{\bm{\theta}}$ is a real algebraic variety defined by the polynomial equation $f_{\bm{\theta},c}(\bm{x})-f_{\bm{\theta},c^{\prime}}(\bm{x})=0$. Imposing the additional inequality constraints $f_{\bm{\theta},i}(\bm{x})-f_{\bm{\theta},c}(\bm{x})\leq 0$ for all $i\neq c,c^{\prime}$ restricts this to a semialgebraic set $\mathcal{B}_{c,c^{\prime}}^{\bm{\theta}}$.
@@
$\mathcal{L}(\bm{x},\lambda,\bm{\mu})$={} $\frac{1}{2}\,d_{\bm{\xi}}(\bm{x})+\lambda\bigl(f_{\bm{\theta},c^{\prime}}(\bm{x})-f_{\bm{\theta},c}(\bm{x})\bigr)+\sum_{i\neq c,c^{\prime}}\mu_{i}\bigl(f_{\bm{\theta},i}(\bm{x})-f_{\bm{\theta},c}(\bm{x})\bigr),$
@@
Since all equations are polynomial, the KKT system can be solved using tools from numerical algebraic geometry. We consider specifically homotopy continuation because it is a complete numerical method guaranteed to find all isolated complex solutions (Li, 1997; Sommese et al., 2001; Sommese and Wampler, 2005; Sommese et al., 2005; Duff et al., 2018). This computes all complex solutions to the stationarity, complementary slackness, and equality constraints. Subsequently, one retains the minimal real solution satisfying the remaining feasibility conditions. Fig.˜2 illustrates the concept, with details in Section˜B.
@@
The proof is provided in Section˜A.1. As a consequence, we may consider the relaxed optimization problem
@@
These yield a square system of $n+1$ equations in $n+1$ unknowns, instead of n+k1n+k-1, resulting in substantial computational savings when the number of classes $k$ is large. This relaxed polynomial system forms the basis of our exact certification procedure via numerical homotopy continuation, which is summarized in Algorithm 1 and described in more detail in  Section˜C.
Algorithm 1 Exact robustness certification via homotopy continuation
1:Network parameters $\bm{\theta}$, test point $\bm{\xi}\in\mathbb{R}^{n}$, radius $\epsilon>0$
2:Relaxed robustness margin $\tilde{\gamma}$ and certification decision
[…]
5:for each $c^{\prime}\in[k]\setminus\{c\}$ do
6:  Define the relaxed critical system
$F_{c,c^{\prime}}^{\bm{\theta}}(\bm{x},\lambda;\bm{\xi}):=\begin{bmatrix}\bm{x}-\bm{\xi}+\lambda\bigl(\nabla f_{\bm{\theta},c^{\prime}}(\bm{x})-\nabla f_{\bm{\theta},c}(\bm{x})\bigr)\\[5.69054pt] f_{\bm{\theta},c^{\prime}}(\bm{x})-f_{\bm{\theta},c}(\bm{x})\end{bmatrix}$.
7:  Construct a homotopy
$H_{c,c^{\prime}}(\bm{z},t;\gamma)=\gamma(1-t)G(\bm{z})+tF_{c,c^{\prime}}^{\bm{\theta}}(\bm{z};\bm{\xi}),\qquad\bm{z}=(\bm{x},\lambda),$
8:  where $G$ is a start system with known solutions and $\gamma\in\mathbb{C}^{\ast}$ is generic
9:  Track all solution paths from $t=0$ to $t=1$
10:  Let $\mathcal{R}_{c,c^{\prime}}$ be the real endpoints of $F_{c,c^{\prime}}^{\bm{\theta}}(\bm{x},\lambda;\bm{\xi})=0$
11:  if $\mathcal{R}_{c,c^{\prime}}\neq\varnothing$ then
12:   $\tilde{\gamma}\leftarrow\min\!\Bigl\{\tilde{\gamma},\;\min_{(\bm{x},\lambda)\in\mathcal{R}_{c,c^{\prime}}}\sqrt{d_{\bm{\xi}}(\bm{x})}\Bigr\}$
13:  end if
14:end for
15:return certified if $\tilde{\gamma}>\epsilon$, else not certified
2.3 Polynomial Neural Networks
In our theoretical analysis we will consider classifiers that are implemented as polynomial neural networks, defined as follows.
For given $L\geq 1$ and $\bm{h}=(h_{0},h_{1},\dots,h_{L})\in\mathbb{N}_{\geq 0}^{L+1}$, a fully connected feedforward neural network with architecture $\bm{h}$ is a map $f_{\bm{\theta}}:\mathbb{R}^{h_{0}}\to\mathbb{R}^{h_{L}}$ of the form
$f_{\bm{\theta}}=f_{L}\circ\sigma_{L-1}\circ f_{L-1}\circ\cdots\circ\sigma_{1}\circ f_{1},$
where, for each $\ell\in[L]$,
$f_{\ell}(\bm{x})=W_{\ell}\bm{x}+\bm{b}_{\ell},\qquad W_{\ell}\in\mathbb{R}^{h_{\ell}\times h_{\ell-1}},\quad\bm{b}_{\ell}\in\mathbb{R}^{h_{\ell}},$
and $\sigma_{\ell}:\mathbb{R}^{h_{\ell}}\to\mathbb{R}^{h_{\ell}}$ is an activation function that acts coordinatewise, i.e.,
$\sigma_{\ell}(x_{1},\dots,x_{d_{\ell}})=\bigl(\phi_{\ell}(x_{1}),\dots,\phi_{\ell}(x_{d_{\ell}})\bigr),$
for some scalar activation function $\phi_{\ell}:\mathbb{R}\to\mathbb{R}$. The parameter of the network is the tuple $\bm{\theta}=(\left(W_{\ell},\bm{b}_{\ell}\right))_{\ell=1}^{L}$ consisting of the weights $W_{\ell}$ and biases $\bm{b}_{\ell}$ of each layer $\ell\in[L]$. The network is shallow if $L=2$ and deep if $L>2$. In this paper, we focus on polynomial neural networks, which have activation functions $\phi_{\ell}$ that are polynomial. More specifically, we consider a monomial activation fixed across layers,
$\phi_{\ell}(t)=t^{d},\qquad\ell\in[L-1],$
for some fixed integer $d\geq 1$. Equivalently, the coordinatewise activation maps are
$\sigma_{\ell}(x_{1},\dots,x_{d_{\ell}})=(x_{1}^{d},\dots,x_{d_{\ell}}^{d}),\qquad\ell\in[L-1].$
In this setting, each component of the network map $f_{\bm{\theta}}$ is a polynomial in the input $\bm{x}$ and also a polynomial in the parameter $\bm{\theta}$}. If all biases $\bm{b}_{\ell}$ are fixed to zero, then the resulting network map is homogeneous. For $k$-class classification, we take $h_{0}=n$ and $h_{L}=k$, so $f_{\bm{\theta}}:\mathbb{R}^{n}\to\mathbb{R}^{k}$ has output coordinates $f_{\bm{\theta},1},\dots,f_{\bm{\theta},k}$ representing the class logits, and the predicted class is $\arg\max_{i\in[k]}f_{\bm{\theta},i}(\bm{x})$.
@@ 3 ED Degree of a Decision Boundary
The ED degree of $\mathcal{V}_{c,c^{\prime}}^{\bm{\theta}}$ is an upper bound on the optimization complexity: for our data point $\bm{\xi}$, the number of complex critical points is bounded above by $\mathrm{EDdegree}(\mathcal{V}_{c,c^{\prime}}^{\bm{\theta}})$, with equality holding when $\bm{\xi}$ is generic.